OpenStudy (anonymous):

How will adding the value 8 affect the mean and median of the data set 4, 6, 8, 10, 12? A. The mean and median will increase by the same amount. B. The median will increase more than the mean. C. The median will stay the same, but the mean will increase. D. The mean and median will both stay the same.

OpenStudy (anonymous):

@jtbulin

OpenStudy (kc_kennylau):

Calculate the (old and new)(mean and median)

OpenStudy (anonymous):

huh?

OpenStudy (kc_kennylau):

mean = sum of all values / number of values median = the middle value or the two middle values/2

OpenStudy (kc_kennylau):

Calculate the old mean, old median, new mean and new median

OpenStudy (anonymous):

sorry hun. \[4+6+8+10+12=40\] \[40/5=8\] mean=8 if you add 8, then it is \[4+6+8+8+10+12=48\] \[48/6= 8.\] the mean is the same. Now for the median.

OpenStudy (anonymous):

 if you add 8 then it is:  again, they stay the same. its A.

OpenStudy (kc_kennylau):

Or it's D because the question is not well set

OpenStudy (the_fizicx99):

Erm, no. Not A. The median is 4,6, *8*, 10 and 12 The median is 8, without 8, it would be the 2 middle terms (10 + 6) / 2 = 16 / 2 = 8 They both stay the same. It's D*

OpenStudy (anonymous):

oh sorry, that was wrong. it is d. Sorry

OpenStudy (anonymous):

it is k

OpenStudy (kc_kennylau):

@tHe_FiZiCx99 the old set is 4,6,8,10,12 and the new set is 4,6,8,8,10,12

OpenStudy (the_fizicx99):

I think they're implying adding the value 8 into the set? One was without the 8 and the other is with the 8?

OpenStudy (anonymous):

@kc_kennylau has it right. They would make you add it, they would not already have it there in the second set you have two 8's

OpenStudy (the_fizicx99):

Well, it says to add, there's already an 8, so it would imply add another 8. Sorry.

OpenStudy (anonymous):

right. you would have two of them

OpenStudy (kc_kennylau):

@tHe_FiZiCx99 your privacy setting won't allow me to reply to your message because I'm not your fan

OpenStudy (the_fizicx99):

I don't have them on?

OpenStudy (kc_kennylau):

Oh I can reply now never mind